OpenStudy (anonymous):

Carlos, Jackie, Lester, and Margie ate lunch at a restaurant. The total amount of the bill, inculding tax and tip, was $44.60. Carlos paid $15.00, Jackie payed 1/4 of the bill, Lester payed 20% of the bill, and Margie payed the rest of the bill. Who paid the greatest part of the bill?

OpenStudy (anonymous):

Okay, so your learning fraction, percentages, and proportions. So let convert them so we can compare the dollar values :) 15$, 25% , and 20% so lets convert these percentages to dollar bills, to do this multiply .25 x 44.60= $ .2 x 44.60 = $

OpenStudy (anonymous):

@Akosua_Boateng ??? Do you need a deeper explanation?

OpenStudy (anonymous):

Let c, j, L and m be Carlos, Jackie, Lester, and Margie respectively. Solve the following equatons for c, j, L and M. { c=15, j=44.60/4, L=0.20*44.60, m=44. 60-c-j-L }

OpenStudy (anonymous):

Okay so we get 15$ by Carlos $11.15 by Jackie $8.92 by Lester The Rest by Margie so add them all together then subtract to find margie at $9.53, so Carlos paid the most :) If you understand the equation posted above than that even more awesome :D

OpenStudy (anonymous):

@Akosua_Boateng Any q's?

OpenStudy (anonymous):

thanks you tons!

OpenStudy (anonymous):

Your welcome :) remember to find any percentage of a number just convert to a decimal and multiply to that number. To get the decimal from percentage, move 2 decimal places to the left ex. 100% = .1 and 50% = .5 :D

OpenStudy (anonymous):

how would i turn like carlos amount into money

OpenStudy (anonymous):

Carlos paid 15$, do mean like a percentage? what do u mean into money?

OpenStudy (anonymous):

oh wait just kidding. i mean like 1/4 how would i turn that into money

OpenStudy (anonymous):

Oh okay :) Well I took the fraction 1/4, and turned it into a percentage 25% or .25 to find how much 1/4 or 25% is of 44.6 in $'s. So multiply .25 by 44.6 to get 11.15$ paid by Jackie. Need a better explanation ?

OpenStudy (anonymous):

thank you alright hopefully i can remember this for my dca on tuesday oh man

OpenStudy (anonymous):

Right, I've been helping my little bro with converting fractions into decimals, so just remember the things i mentioned above and you'll do great :) Take the percentage convert to a decimal and multiply :) 3/5 60% .6 to get 60% or 3/5 of a number like 30 just multiply .6 x 30 = 18 :)

OpenStudy (anonymous):

Any other q's :)

OpenStudy (anonymous):

nope thank you for the help i really needed it :)

OpenStudy (anonymous):

Awesome, and good luck! happy to help :)